Use Substitution to solve each system of equations. If the system does not have exactly one solution, state whether it has no solution or infinitely many solutions. 18) 9x^2-6x+1
Answer by checkley77(12844)

x-3y=-4 multiply by 2 & add.
2x+6y=5
2x-6y=-8
-----------------
4x=-3
x=-3/4 ans.
-3/4-3y=-4
-3y=-4+3/4
-3y=(-4*4+3)/4
-3y=(-16+3)/4
-3y=-13/4
-3y*4=-13
-12y=-13
y=-13/-12
y=13/12 ans.